OpenStudy (anonymous):

Minimize S = x + 8y with xy = 8 and both x and y > 0.

OpenStudy (raden):

xy=8 ---> y=8/x S = x + 8y S = x + 8(8/x) S = x + 64/x take S' = 0, solve for x!

OpenStudy (raden):

can u find derivative of S

OpenStudy (anonymous):

Do you kno what the answer is ?

OpenStudy (raden):

i got x=8 and y=1 to minimize of S so, S=x + 8y=8 + 8(1) = 8 + 8 = 16

OpenStudy (anonymous):

Yeah same thing weith me thanks for the help